# TopLoc: A Locality Sensitive Hashing Scheme for Trustless Verifiable Inference

### v1

@@ 5.1 Dataset and Models
For our experiments, we use the UltraChat dataset (Ding et al., 2023). The UltraChat dataset contains 1.4 million dialogues consisting of real-world inquiries, creative writing prompts, and various other text-based tasks such as rewriting, continuation, summarization and inference covering a wide range of topics.
We conduct experiments with three models: Llama-3.1-8B-instruct (Grattafiori et al., 2024), INTELLECT-1-instruct (Jaghouar et al., 2024), and Gemma-2-9b-it (Team et al., 2024), aiming to capture diversity across different architectural dimensions. Llama-3.1-8B-instruct and Intellect-1-instruct share a similar transformer block architecture but differ in the number of layers, while Gemma-2-9b-it features a different hidden dimension and MLP activation function.
@@ 5.2 Experiment Setup
We use bf16 precision for all our experiments unless otherwise specified as bf16 is the most unstable floating point representation commonly used in practice for activations in language model inference, only containing 7 bits of mantissa as opposed to the 23 mantissa bits of fp32 and 10 bit mantissa of fp16.
For all our experiments, we perform the generation autoregressively, generating each new token separately with KV-caching. For the validation we use a prefill to obtain the last hidden state activations for all tokens at once. This allows the validation to be done up to 100×100$\times$100 × (Agrawal et al., 2024, 2023) faster than the generation as the prefill is significantly more compute intensive than the autoregressive decodes and is thus able to better utilize GPU resources.
Figure 2: When generating the response, we need to perform one prefill for the input tokens and then multiple decodes for each new token generated. When validating, we can pass all the tokens at once and perform just one prefill. Decodes are not efficient on GPUs because they are memory bound. Notice that the sequential nature of generation causes us to need the decode blocks multiple times in the generation computation. This increases the total amount of data movement required to perform the computation. The decodes are thus bottlenecked by the time it takes to move data from GPU HBM (High Bandwidth Memory) to SRAM (Shared Memory).
For the thresholds, we use $T_{exp}=90$, $T_{mean}=10$ and $T_{median}=8$ for bf16 inference and $T_{exp}=120$, $T_{mean}=256$ and $T_{median}=128$ for fp32 inference. These thresholds were picked based on our results from analyzing the error statistics in Table 9 and Table 5.
5.3 High Magnitude Activations Exhibit Low Error Rates
As we wish to distinguish inferences using the top-k values in the activations, a key assumption of our method is that high-magnitude elements in the activations are less prone to errors. In Section 3.3, we present the theoretical basis for this hypothesis. In this section, we collect experimental evidence supporting it.
Table 1 presents the error in exponent bits for the 2048th decoded token across various top-k values from 2000 sampled queries using the Llama-3.1-8B-Instruct model. The results highlight the relationship between activation magnitude and error rate, notably that the magnitude of errors generally increases with higher values of top-k.
Table 1: Exponent bit error counts for the 2048th decoded token across various top-k values in 2000 queries using Llama-3.1-8B-Instruct. The results illustrate that exponent errors increase in magnitude as the number of top-k values we take from the tensor increases. Deviations with magnitude above 100 are a result of catastrophic cancellations and mostly appear in the bottom 50% of values in the tensor.
@@ 5.3 High Magnitude Activations Exhibit Low Error Rates
We also plot the maximum percentage of elements with exponent errors from our experiment on 2000 generations in Figure 3. Lower values of top-k have a lower percentage of exponent error in general.
Figure 3: Maximum exponent bit error rate as a function of top-k elements across different models for 2000 queries. This demonstrates that exponent bit errors increase consistently as more elements are included in the top-k set across various model architectures.
@@ 5.4 Deviations in the Mantissa Are Small When the Exponent Bits Are Matched
In floating point computations, deviations in the mantissa are often amplified when the exponent bits are incorrect. However, when the exponent bits are matched, the observed deviations in the mantissa remain relatively small, even as the number of decode tokens is increased.
We analyzed the absolute differences in the mantissa for the top 128 elements of the last hidden layer activations across 2000 queries. The results indicate that while there is a noticeable increase in mantissa errors as the token index grows, the overall growth rate is moderate. This suggests that the impact of token index on floating-point precision is not severe, even for higher indices.
To investigate this, we analyzed the absolute differences in the mantissa for the top 128 elements of the last hidden layer activations across 2,000 queries. Our results, visualized in Figure 4, reveal a noticeable but moderate increase in mantissa errors as the token index grows. This suggests that, while longer generations do introduce some additional mantissa errors, the overall impact on floating-point precision remains limited.
This shows that the mantissa can serve as a useful indicator for validating legitimate computations. Even with the inherent randomness introduced by hardware and algebraic reordering, the small deviations in the mantissa when the exponent bits are correct suggest that the mantissa can effectively distinguish between expected computational outcomes and potential attempts at foul play.
These findings suggest the potential of using the mantissa error mean and median statistics as a robust indicator for validating legitimate computations. Despite the randomness introduced by hardware variability and algebraic reordering, the relatively small deviations observed in the mantissa (when the exponent bits are correct) demonstrate its effectiveness in distinguishing between expected computational outcomes and potential anomalies or attempts at manipulation.
Figure 4: Impact of token index on mantissa errors in the top 128 elements of the last hidden activations across 2000 queries. The results demonstrate that while mantissa errors increase with higher token indices, the growth rate remains moderate, indicating a limited impact of token index on floating-point precision.
@@ 5.5 Mismatch Rate for Different Values of Top-k
An issue with comparing top-k values is that the top-k values may not be the same in the tensors being compared.
Figure 5 illustrates the mismatch error ratio for top-k elements across different models, highlighting how the amount of top-k matches varies as the value of top-k increases. The results demonstrate that the mismatch error decreases significantly with larger values of top-k. This is because the boundary between elements that are in the top-k and those that are narrowly excluded is the source of mismatch. This boundary becomes smaller relative to the set of top-k elements as the size of the set increases, ultimately becoming 0 when all tensor elements are used. For smaller top-k values, the maximum mismatch remains relatively low, suggesting that discrepancies in element alignment are minimal even for small k. Furthermore, the median mismatch is consistently an order of magnitude lower than the maximum mismatch, indicating that most errors are minor and well within acceptable limits.
Figure 5: Max and median mismatch error ratio for top-k elements across different models.
These findings reveal that the top-k indexes can be reproduced reliably, even in the presence of numerical variations.
@@ 5.6 Robustness across Different GPUs, Attention and Tensor Parallel Implementations
To assess robustness across tensor parallel configurations and GPU setups, we run the inference using vLLM [^4] https://github.com/vllm-project/vllm (Kwon et al., 2023) with a hook to obtain the top-k values from the last hidden layer. We thus use the vLLM implementation of tensor parallel and paged attention. We generate 512 new tokens for 400 prompts. In order to reduce the amount of values we need to store, we save the top-128 values every 32 new tokens. This results in $512/32=16$ sets of top-128 values for the decode activations. We also save a set of top-128 values for the activations computed for the input prompt. The experiments were performed on 3 models, Llama-3.1-8B-Instruct, Intellect-1-Instruct and Gemma-2-9b-it which are then aggregated.
Table 2 presents the success rates for different combinations of tensor parallel configurations and GPUs. As shown, TopLoc achieves a 100% success rate across all tested configurations, including combinations where the generation model and validation model utilized different hardware. We also tabulate the worst case error statistics for the validations in Table 3. Here it is clear that none of the error statistics exceed the thresholds proposed in Section 5.2.
Table 2: Validation success rate with different tensor parallel configurations and GPUs
Validation Model
Generation Model
1 x A100 1 x 4090 2 x 4090
1 x A100 $100\%$ $100\%$ $100\%$
1 x 4090 $100\%$ $100\%$ $100\%$
2 x 4090 $100\%$ $100\%$ $100\%$
Table 3: Error statistics for validation with different tensor parallel configurations and GPUs
@@
Min Top-k
Intersection
Min Exponent
Match
@@
1xA100 1xA100 $118\;(92.19\%)$ $112\;(87.50\%)$ $5.06$ $2$
1x4090 $118\;(92.19\%)$ $110\;(85.94\%)$ $4.68$ $2$
2x4090 $113\;(88.28\%)$ $109\;(85.16\%)$ $4.96$ $4$
1x4090 1xA100 $119\;(92.97\%)$ $112\;(87.50\%)$ $6.30$ $3$
1x4090 $119\;(92.97\%)$ $108\;(84.38\%)$ $4.03$ $3$
2x4090 $119\;(92.97\%)$ $108\;(84.38\%)$ $4.03$ $3$
2x4090 1xA100 $117\;(91.41\%)$ $113\;(88.28\%)$ $5.15$ $2$
1x4090 $117\;(91.41\%)$ $113\;(88.28\%)$ $5.16$ $4$
2x4090 $116\;(90.62\%)$ $113\;(88.28\%)$ $5.15$ $3$
@@
We further evaluated TopLoc’s robustness by testing different attention kernel implementations for the generation and validation. We use Hugging Face transformers and its references to the attention implementations, FlashAttention v2 (Flash2), PyTorch Scaled Dot Product Attention (SDPA), and FlexAttention (Flex).
Table 8 summarizes the success rates achieved across these configurations. Similar to the GPU experiments, TopLoc exhibited a 100% success rate for all combinations of generation and validation models using different attention kernel implementations. We also tabulate the worst case error statistics for the validations in Table 5.
Table 4: Validation success rate with different attention kernels.
Validation Model
Generation Model
Flash2 SDPA Flex
Flash2 $100\%$ $100\%$ $100\%$
SDPA $100\%$ $100\%$ $100\%$
Flex $100\%$ $100\%$ $100\%$
Table 5: Error statistics with different attention kernel combinations.
Generation
Model
Validation
Model
Min Top-k
Intersection
Min Exponent
Match
Max Mantissa
Diff Mean
Max Mantissa
Diff Median
Flash Flash $103\;(80.47\%)$ $100\;(78.12\%)$ $4.64$ $2$
SDPA $117\;(91.41\%)$ $112\;(87.50\%)$ $4.25$ $2$
Flex $103\;(80.47\%)$ $100\;(78.12\%)$ $3.79$ $3$
SDPA Flash $120\;(93.75\%)$ $110\;(85.94\%)$ $3.85$ $3$
SDPA $119\;(92.97\%)$ $114\;(89.06\%)$ $3.15$ $3$
Flex $117\;(91.41\%)$ $111\;(86.72\%)$ $3.88$ $3$
Flex Flash $118\;(92.19\%)$ $112\;(87.50\%)$ $3.72$ $2$
SDPA $119\;(92.97\%)$ $112\;(87.50\%)$ $4.02$ $2$
Flex $121\;(94.53\%)$ $116\;(90.62\%)$ $3.43$ $2$
5.7 Distinguishes Models, Prompts and Compute Precision
We conducted experiments to assess the ability of TopLoc to detect variations in models, alterations to the input prompt and use of different precision. The results demonstrate the method’s reliability in distinguishing these changes.
To asses the ability of TopLoc to differentiate between models, we generate commits using four different models, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, Intellect-1-Instruct and Gemma-2-9b-it. We then validate the generations using the four models and tabulate the success rate of validation in Table 6.
Table 6: Success rate of TopLoc in differentiating between generation and validation model computations using 2,000 prompts from the UltraChat dataset (2,048 new tokens sampled during decode per prompt). When the generation and validation models are the same, the success rate is 100%, indicating no false negatives. When the models are different, the success rate is 0%, demonstrating that our method has no false positives and effectively distinguishes between models.
Validation Model
Generation Model
Llama-3.1
8B-Instruct
Llama-3.1
70B-Instruct
Intellect-1
10B-Instruct
Gemma-2
9b-it
Llama-3.1-8B-Instruct $100\%$ $0\%$ $0\%$ $0\%$
Llama-3.1-70B-Instruct $0\%$ $100\%$ $0\%$ $0\%$
Intellect-1-Instruct $0\%$ $0\%$ $100\%$ $0\%$
Gemma-2-9b-it $0\%$ $0\%$ $0\%$ $100\%$
We further evaluated TopLoc’s robustness by testing it on three types of altered prompts:
@@ 5.7 Distinguishes Models, Prompts and Compute Precision
  • Advertising: A system prompt that asks the model to advertise a vitamin supplement when asked about health.
@@
  • Taco: A short prompt that directs the model to always praise tacos.
The full prompts can be found in Appendix B.3.
The success rates for differentiating models with these altered prompts are summarized in Table 7. For all tested scenarios, TopLoc correctly identifies when the input prompt has been altered.
Table 7: Validation success rate with altered prompt
Model
Advertising Avoidance Taco prefix
Llama-3.1-8B-Instruct $0\%$ $0\%$ $0\%$
Intellect-1-Instruct $0\%$ $0\%$ $0\%$
Gemma-2-9b-it $0\%$ $0\%$ $0\%$
We also tested for the ability to differentiate models based on compute precision. Specifically, we evaluated the success rate when using 32-bit floating point (fp32) and 16-bit brain floating point (bf16) computations. The results in Table 8 demonstrate TopLoc’s robustness, with a 100% success rate in identifying matching precisions and distinguishing mismatched ones.
Table 8: Success rate with different precision
Validation Model
Generation Model
fp32 bf16
fp32 $100\%$ $100\%$
bf16 $0\%$ $100\%$
To further understand the acceptance boundaries, we measured error metrics such as top-k intersection, exponent match, and mantissa differences across decode-prefill precision combinations. Table 9 summarizes these comparisons, providing insights into the numerical effects of precision mismatches.
Due to the differing number of mantissa bits between fp32 and bf16 formats, we scaled the fp32 mantissa to match bf16 when validating with bf16. Conversely, when validating a bf16 decode model with fp32, we padded 16 zero bits to the bf16 representation. These adjustments ensured fair comparisons despite the inherent differences in precision formats.
Table 9: Error comparisons of different Generation-Validation precision combinations. The upward arrow ($\uparrow$) indicates a maximum value, while the downward arrow ($\downarrow$) indicates a minimum value.
@@
fp32 fp32 $127\downarrow\;(99.22\%)$ $127\downarrow\;(99.22\%)$ $180.38\uparrow$ $48\uparrow$
bf16 $122\downarrow\;(95.31\%)$ $114\downarrow\;(89.06\%)$ $4.38\uparrow$ $2\uparrow$
bf16 fp32 $128\uparrow\;(100\%)$ $128\uparrow\;(100\%)$ $27892.02\downarrow$ $21683\downarrow$
bf16 $119\downarrow\;(92.97\%)$ $112\downarrow\;(87.50\%)$ $6.30\uparrow$ $3\uparrow$

### v2

@@ 5.1 Dataset and Models
For our experiments, we use the UltraChat dataset (Ding et al., 2023). The UltraChat dataset contains 1.4 million dialogues consisting of real-world inquiries, creative writing prompts, and various other text-based tasks such as rewriting, continuation, summarization, and inference, covering a wide range of topics.
We conduct experiments with three models: Llama 3.1-8B-instruct (Grattafiori et al., 2024), INTELLECT-1-instruct (Jaghouar et al., 2024), and Gemma-2-9b-it (Gemma Team et al., 2024), aiming to capture diversity across different architectural dimensions. Llama-3.1-8B-instruct and Intellect-1-instruct share a similar transformer block architecture but differ in the number of layers, while Gemma-2-9b-it features a different hidden dimension and MLP activation function.
@@ 5.2 Experiment Setup
We use the bf16 precision for all our experiments, unless specified otherwise. bf16 is commonly used in practice for activations in language model inference. However, compared to fp16 and fp32 precision, it is most prone to catastrophic cancellations. It contains only 7 bits of mantissa, as opposed to 23 mantissa bits for fp32 and 10 bits for fp16.
For all of our experiments, we perform the generation autoregressively, generating each new token separately with KV caching. During validation, we obtain the last hidden state activations for all tokens at once. This allows the validation to be done significantly faster than the generation, as the prefill is significantly more compute-intensive than the autoregressive decoding and is thus able to better utilize the GPU resources (Agrawal et al., 2024, 2023).
For thresholds, we use $T_{exp}=38$, $T_{mean}=10$ and $T_{median}=8$ for bf16 inference and $T_{exp}=8$, $T_{mean}=256$ and $T_{median}=128$ for fp32 inference. These thresholds were chosen based on our analysis of the error statistics in Table 2 and Table 5.
5.3 High-Magnitude Activations Have Low Error Rates
Table 1: Exponent bit error counts for the 2048th decoded token across various top-k values in 2000 queries using Llama-3.1-8B-Instruct.
Top-k Exact Match Small Deviations Larger Deviations
@@ 5.3 High-Magnitude Activations Have Low Error Rates
64 126,973 512 508 4 - 3 - - -
128 254,693 761 529 11 - 5 - - 1
256 502,130 5,824 3,993 38 - 14 - - 1
512 1,002,724 10,693 10,471 80 - 31 - - 1
1024 2,023,123 13,159 11,222 342 2 150 - - 2
2048 3,997,083 49,340 47,661 1,142 41 727 - - 6
4096 7,495,155 296,584 298,716 27,350 27,169 15,569 15,386 - 16,071
As we wish to distinguish inference results using the top-k magnitude elements in the activations, a key assumption of our method is that high-magnitude elements in the activations are less prone to errors. In Section 3.3, we present the theoretical basis for this hypothesis. Here, we collect the experimental evidence supporting it.
Table 1 presents the error in exponent bits for the 2048th decoded token across various top-k values from 2000 sampled queries using Llama-3.1-8B-Instruct. The results highlight the relationship between activation magnitude and error rate, notably that the magnitude of errors generally increases with higher values of top-k. Deviations with a magnitude above 100 are the result of catastrophic cancellations and mostly appear in the bottom 50% of values in the tensor.
@@ 5.4 Deviations in the Mantissa Are Small When the Exponent Bits Are Matched
Figure 2: Impact of token index on mantissa errors in the top 128 elements of the last hidden activations across 2000 queries. The errors increase as the token index grows because later tokens rely more on KV cache values with compounding errors. However, the increase is moderate, suggesting that the errors grow in a limited manner.
In floating point computations, mantissa deviations are often amplified by mismatched exponent bits but remain relatively small when the exponent bits match.
We analyzed the absolute differences in the mantissa for the top 128 elements of the last hidden layer activations across 2,000 queries using Llama 3.1-8B-Instruct. Our results, shown in Figure 2, indicate an increase in mantissa errors as the token index increases. This increase occurs because the errors in the KV cache compound, causing higher token indices to have a higher deviation as the inputs in the forward pass become more dependent on the cached values. However, the increase is moderate, suggesting that longer generations introduce only limited floating-point precision errors, even at higher token indices.
The findings highlight the mantissa as a useful indicator for validating computations. When the exponent bits are matched, the mantissa deviations remain small despite hardware variability and algebraic reordering. This suggests that mantissa error mean and median statistics can effectively detect computational anomalies or attempts at manipulation.
@@ 5.5 Mismatch Rate for Different Values of Top-k
An issue with comparing top-$k$ values is that the top-k indices may not be the same in the tensors being compared.
Figure 3 illustrates the mismatch error ratio for top-$k$ indices across different models. The results demonstrate that the mismatch error decreases significantly with larger values of top-$k$. This is because the boundary between elements that are in the top-$k$ and those that are narrowly excluded is the source of mismatch. This boundary becomes smaller relative to the set of top-k elements as the size of the set increases, ultimately becoming 0 when all tensor elements are used. For smaller top-k values, the maximum mismatch remains relatively low, suggesting that discrepancies in element alignment are minimal even for small $k$. Furthermore, the median mismatch is consistently an order of magnitude lower than the maximum mismatch, indicating that most errors are minor and well within acceptable limits.
Figure 3: Max and median mismatch error ratio for top-k indices across different models. The error ratio is obtained by comparing the top-k indices as a set between the generation and the validation. The ratio decreases as the elements in the top-k set increases, indicating a slower growth in the number of elements slipping past the top-k cutoff.
These findings reveal that the top-$k$ indices can be reproduced reliably, even in the presence of numerical variations.
@@ 5.6 Robustness across Different GPUs, Attention and Tensor Parallel Implementations
Table 2: Error statistics for validation with different tensor parallelism configurations, GPUs and attention kernel implementations. All the error statistics are below the thresholds ($T_{exp}=38$, $T_{mean}=10$, $T_{median}=8$), indicating that they are all classified correctly as positive samples.
@@
Max Top-k
Mismatch
Max Exponent
Mismatch
@@
1xA100 1xA100 $10\;(7.81\%)$ $16\;(12.50\%)$ $5.06$ $2$
1x4090 $10\;(7.81\%)$ $18\;(14.06\%)$ $4.68$ $2$
2x4090 $15\;(11.72\%)$ $19\;(14.84\%)$ $4.96$ $4$
1x4090 1xA100 $9\;(7.03\%)$ $16\;(12.50\%)$ $6.30$ $3$
1x4090 $9\;(7.03\%)$ $20\;(15.62\%)$ $4.03$ $3$
2x4090 $9\;(7.03\%)$ $20\;(15.62\%)$ $4.03$ $3$
2x4090 1xA100 $11\;(8.59\%)$ $15\;(11.72\%)$ $5.15$ $2$
1x4090 $11\;(8.59\%)$ $15\;(11.72\%)$ $5.16$ $4$
2x4090 $12\;(9.38\%)$ $15\;(11.72\%)$ $5.15$ $3$
Flash Flash $25\;(19.53\%)$ $28\;(21.88\%)$ $4.64$ $2$
SDPA $11\;(8.59\%)$ $16\;(12.50\%)$ $4.25$ $2$
Flex $25\;(19.53\%)$ $28\;(21.88\%)$ $3.79$ $3$
SDPA Flash $8\;(6.25\%)$ $18\;(14.06\%)$ $3.85$ $3$
SDPA $9\;(7.03\%)$ $14\;(10.94\%)$ $3.15$ $3$
Flex $11\;(8.59\%)$ $17\;(13.28\%)$ $3.88$ $3$
Flex Flash $10\;(7.81\%)$ $16\;(12.50\%)$ $3.72$ $2$
SDPA $9\;(7.03\%)$ $16\;(12.50\%)$ $4.02$ $2$
Flex $7\;(5.47\%)$ $12\;(9.38\%)$ $3.43$ $2$
To assess robustness across tensor parallel configurations and GPU setups, we run the inference using vLLM (Kwon et al., 2023) with a hook to obtain the top-k values from the last hidden layer. We thus use the vLLM implementation of tensor parallelism and PagedAttention. We generate 512 new tokens for 400 prompts. In order to reduce the amount of values we need to store, we save the top-128 values every 32 new tokens. This results in $512/32=16$ sets of top-128 values for the decode activations. We also save a set of top-128 values for the activations computed for the input prompt. The experiments were performed on 3 models, Llama-3.1-8B-Instruct, Intellect-1-Instruct and Gemma-2-9b-it which are then aggregated.
@@
We further evaluated TopLoc’s robustness by testing different attention kernel implementations for the generation and validation. We use Hugging Face Transformers and its references to the attention implementations, Flash Attention 2, PyTorch Scaled Dot Product Attention, and FlexAttention.
We report the worst-case error statistics for different tensor parallelism and GPU combinations in Table 2. Here, none of the error statistics exceed the thresholds proposed in Section 5.2. Lastly, we display the worst-case error statistics for attention kernel combinations in Table 2. The statistics do not exceed the proposed thresholds in Section 5.2, indicating that all generated proofs would have been accepted.
5.7 TopLoc Distinguishes Models, Prompts and Compute Precision
Table 3: Error statistics for different prompt alterations. The minimum error statistics for each of the prompt alterations are above the exponent mismatch threshold of $38$, indicating that all the samples are correctly classified as negative samples.
Prompt Alteration Min Exponent Mismatch
Advert $95\;(74.22\%)$
Avoid $74\;(57.81\%)$
Taco $67\;(52.34\%)$
Table 4: Error statistics for validation with different models. The upward arrow ($\uparrow$) indicates a maximum value, while the downward arrow ($\downarrow$) indicates a minimum value. Given the thresholds $T_{exp}=38$, $T_{mean}=10$, $T_{median}=8$, the minimum errors for mismatching pairs are all above the threshold and are all correctly classified as negatives. The maximum errors for matching pairs are below the thresholds and are all correctly classified as positives. Mismatch combinations with Gemma-2-9b-it as the generation model are not shown because they will fail the validation by having an out-of-bound token in the completion due to the bigger vocabulary size of Gemma-2-9b-it.
Generation Model Validation Model Top-k
Mismatch
Exponent
Mismatch
Mantissa Diff
Mean/Median
Llama-3.1-8B-Instruct Llama-3.1-8B-Instruct $8\uparrow\;(6.2\%)$ $13\uparrow\;(10.2\%)$ $4.18$/$4\uparrow$
Llama-3.1-70B-Instruct Llama-3.1-70B-Instruct $7\uparrow\;(5.5\%)$ $17\uparrow\;(13.3\%)$ $2.39$/$2\uparrow$
Intellect-1-Instruct Intellect-1-Instruct $8\uparrow\;(6.2\%)$ $8\uparrow\;(6.2\%)$ $2.52$/$2\uparrow$
Gemma-2-9b-it Gemma-2-9b-it $17\uparrow\;(13.3\%)$ $19\uparrow\;(14.8\%)$ $5.38$/$2\uparrow$
Llama-3.1-8B-Instruct Llama-3.1-70B-Instruct $126\downarrow\;(98.44\%)$ $128\downarrow\;(100\%)$- / --
Intellect-1-Instruct $122\downarrow\;(95.31\%)$ $125\downarrow\;(97.66\%)$- / --
Gemma-2-9b-it $125\downarrow\;(97.66\%)$ $126\downarrow\;(98.44\%)$- / --
Llama-3.1-70B-Instruct Llama-3.1-8B-Instruct $127\downarrow\;(99.22\%)$ $128\downarrow\;(100\%)$- / --
Intellect-1-Instruct $127\downarrow\;(99.22\%)$ $127\downarrow\;(99.22\%)$- / --
Gemma-2-9b-it $126\downarrow\;(98.44\%)$ $126\downarrow\;(98.44\%)$- / --
Intellect-1-Instruct Llama-3.1-8B-Instruct $126\downarrow\;(98.44\%)$ $127\downarrow\;(99.22\%)$- / --
Llama-3.1-70B-Instruct $125\downarrow\;(97.66\%)$ $127\downarrow\;(99.22\%)$- / --
Gemma-2-9b-it $126\downarrow\;(98.44\%)$ $126\downarrow\;(98.44\%)$- / --
To assess the ability of TopLoc to differentiate between models, we generate proofs using four models: Llama 3.1-8B-Instruct, Llama 3.1-70B-Instruct, Intellect-1-Instruct and Gemma-2-9b-it. We then validate the generations using the same models and show the results in Table 4. When the models used for generation and validation are the same, the worst-case error statistics are below the threshold, indicating that they would have passed validation. When they are different, the best-case error statistics exceed the threshold, indicating that they would all have failed validation.
We further evaluated TopLoc’s robustness by testing it on three types of altered prompts. Some of the alterations are long, while some are only 4 tokens. To test the method under different prompt lengths, we select prompts that are multiple sentences, one sentence and just 3 words long.
The full prompts can be found in Appendix B.3, and they are split into the following categories:
@@ 5.7 TopLoc Distinguishes Models, Prompts and Compute Precision
  • Advertising: A system prompt that asks the model to advertise a vitamin supplement when asked about health and wellness-related topics.
@@
  • Taco: A short prompt that directs the model to always praise tacos in the response.
Table 3 shows that for all prompt alterations, the best-case exponent mismatches are above the threshold of 38, indicating that they would all have failed the validation.
We also tested TopLoc for the ability to differentiate models based on the compute precision. Specifically, we evaluated the success rate when using 32-bit floating point (fp32) and 16-bit Brain floating point (bf16) computations.
Due to the differing number of mantissa bits between fp32 and bf16 formats, we scaled the fp32 mantissa to match bf16 when validating with bf16. Conversely, when validating a bf16 decode model with fp32, we padded 16 zero bits to the bf16 representation. These adjustments ensured fair comparisons despite the inherent differences in precision.
Table 5 contains the errors for the different combinations, showing that fp32 is able to pass validations with either precision, while bf16 is only able to pass when the validator is also bf16, always failing if the validator uses fp32.
Table 5: Error comparisons of different Generation-Validation precision combinations. The upward arrow ($\uparrow$) indicates a maximum value, while the downward arrow ($\downarrow$) indicates a minimum value. When the generation model is fp32, all the samples are positive and the maximum errors are all below the thresholds ($T_{exp}=38$, $T_{mean}=10$, $T_{median}=8$). When the generation model is bf16, only the samples where we validate with the bf16 model and threshold are positive. When we validate with the fp32 model and thresholds, the minimum mantissa differences are above the thresholds, indicating that all of the samples are correctly classified as negative.
@@
fp32 fp32 $1\uparrow\;(0.78\%)$ $1\uparrow\;(0.78\%)$ $180.38\uparrow$ $48\uparrow$
bf16 $6\uparrow\;(4.69\%)$ $14\uparrow\;(10.94\%)$ $4.38\uparrow$ $2\uparrow$
[…]
bf16 $9\uparrow\;(7.03\%)$ $16\uparrow\;(12.50\%)$ $6.30\uparrow$ $3\uparrow$
5.8 Overhead and detection rate compared to prior methods and baselines
Table 6: Time and memory overhead of generating and validating Llama-2-13B inference using established verifiable inference compared to our method. TopLoc is way cheaper and significantly more practical compared to prior approaches. The memory overhead is millions of times lower compared to zkLLM and $98,000$x less compared to SVIP. TopLoc can also be used out of the box, unlike SVIP which requires training a detection model for each model we wish to detect. TopLocis also more reliable, having no false positive or false negative rate in the settings we tested.
zkLLM SVIP Raw Activations TopLoc
Commitment size per token 11 MB 20KB 10KB 8B
Committing overhead per token $986s$ $1.7ms$ $0.26ms$
Validation time $803s$ $5.6ms$ $81ms$ $81ms$
Provider GPU memory overhead per token 23.1GB 980MB 10KB 10KB
False Positive Rate 0% 3% 0% 0%
False Negative Rate (Deterministic) 0% 4.41 % 0% 0%
False Negative Rate (Non-Deterministic) 100% - 0% 0%
To evaluate the practical advantages of TopLoc, we compare its key performance metrics such as computational overhead, proof size, and detection accuracy against established verifiable inference techniques like zkLLM (Sun et al., 2024a) and SVIP (Sun et al., 2024b). We also include a baseline where we directly store and validate the raw intermediate activations.
Table 6 provides a summary of these comparisons, underscoring TopLoc’s superior efficiency and its robustness to non-deterministic GPU operations. Compared to previous methods in the literature such as zkLLM and SVIP, TopLoc has significantly reduced memory overhead. TopLoc is also more reliable than prior methods and can be used out of the box for any model, unlike SVIP which requires the training of a detection model.
To quantify the storage efficiency of TopLoc, we compare its memory footprint against a baseline of storing the final hidden activations directly. We illustrate this using the smallest model we tested: Llama-3.1-8B-Instruct model, which has a hidden size of $4096$. Storing the complete final hidden activations for every token generated in bf16 precision would require $2$ bytes per element for $4096$ elements for a total of $8192$ bytes per token. In contrast, TopLoc stores the top-128 activation values, sampled every N=32 tokens, using a polynomial congruence represented by 128 coefficients. With each coefficient requiring 2 bytes, the total storage is 256 bytes per 32-token interval. This amortizes to 8 bytes per token, representing a 1024×1024$\times$1024 × reduction compared to direct storage.